OpenStudy (anonymous):

5/8x + 1/16x = 9/16 + x , x=

OpenStudy (anonymous):

If we see that 5/8 = 10/16 and x = 16/16x we can go 10/16x +1/16x = 9/16 + 16/16x and then drop all the 16's on the bottom 10x + x = 9 + 16x or -5x = 9 and x= -9/5